OpenStudy (anonymous):

create a word problem for systems of equations. School A and School B are selling tickets for their school play.Whoever sells the most tickets gets a mention in the local newspaper.School A sells 40 student tickets and 20 senior tickets for a total of $160.School B sold 10 student tickets and 60 senior tickets for a total of $100.

OpenStudy (anonymous):

is this a good word problem to solve for graphing elimination and substituition. @freckles @phi @Hero

OpenStudy (anonymous):

@ganeshie8

OpenStudy (phi):

not the way it is worded.

OpenStudy (anonymous):

i dont understand what you mean what do i need to change

OpenStudy (phi):

First, the question: who sells the most tickets does not use a "system of equations" to solve you see school A sold 60 and B sold 70. Second, system of equations (with 2 equations) should have two unknowns I don't know what the unknowns are. Maybe price of each type of ticket, but unless both schools charge the same amount for each type (and we can't assume they do, unless you tell us), we are in trouble.

OpenStudy (phi):

You can "tweak" the problem, by saying it's one school, and two people different people sold the tickets. (that way we know both used the same price for the "student ticket" and the "senior ticket"), and the problem is to figure out the price of each type of ticket.

OpenStudy (anonymous):

sooo Bridgette and ana are competing with each other to see who can sell the most tickets for their school play.Bridgette sold 20 senior tickets and 40 student tickets for a total of 110.Ana sold 40 senior tickets and 10 student tickets for a total of 100. find the cost of one senior ticket and one student ticket

OpenStudy (anonymous):

is that better

OpenStudy (phi):

yes, but now you have to solve the system to make sure the numbers make sense. You don't want the price to be negative (for example) The easier way is to let the price be 1.00 and 2.00 then use those numbers to find: 20*1 + 40*2 = 20+80=100 so make the total dollars be 100 and for the other person 40*1 + 10*2= 40+20= 60 . so use 60 for the dollars the second person got.

OpenStudy (phi):

or use your own prices (not mine)

OpenStudy (anonymous):

so you want me to say that one total is 1.00 and the other total is 2.00

OpenStudy (phi):

no, I am saying pick the "answers" (the prices) and use those numbers to figure out the total dollars (in your question)

OpenStudy (phi):

if you pick the prices (I just gave an example using simple numbers like 1 and 2) then you can figure out how much each person got. Am I unclear?

OpenStudy (anonymous):

im kinda confused so u want me to change the amount of tickets

OpenStudy (phi):

If you make up a problem, you should know the answer to the problem (just in case you have to solve it, or check if someone else got the correct answer). when you say Bridgette sold 20 senior tickets and 40 student tickets for a total of 110. how did you pick the 110 ?

OpenStudy (phi):

Instead of randomly picking 110, I am saying "randomly pick the price of a senior ticket and the price of a student ticket" (those are the numbers someone would have to figure out) then using your prices, you do 20*price of senior ticket + 40 * price of student ticket and get a number (not 110) then use those same prices for the other person and get their total. use those totals instead of the 110 and the 100 that you now are using. that way, you know what the answer is (the prices you picked)

OpenStudy (anonymous):

so im supposed to say Bridgette sold 20 senior tickets and 40 student tickets

OpenStudy (phi):

for a total of ****, and you need to put in some dollar amount. I'm saying that you should not just pick 110, but instead, pick the prices and figure out how much the total comes to, using 20 senior and 40 student. can you pick a price for a senior ticket and for a student ticket?

OpenStudy (anonymous):

for a total of 2.00 bridgette sold 20 senior tickets and 40 student tickets.

OpenStudy (phi):

you don't give the answer in the problem. But you should pick the price for a senior ticket. what price do you want to use ?

OpenStudy (anonymous):

6.00

OpenStudy (anonymous):

bridgette is selling senior tickets for 6.00 and student tickets for 3.00 dollars

OpenStudy (phi):

if bridgette sold 20 senior tickets at 6.00 per ticket, how many dollars did she get?

OpenStudy (anonymous):

120

OpenStudy (phi):

and she sold 40 student tickets at 3.00 per ticket. how much is that ?

OpenStudy (anonymous):

120

OpenStudy (phi):

and in total 120+120= 240 so use that number in your problem Bridgette sold 20 senior tickets and 40 student tickets for a total of $240.00 now do the same for the other person. using 6 and 3, figure out how much their total is.

OpenStudy (anonymous):

ana sold 10 senior tickets and 80 student tickets for a total of 240.00

OpenStudy (anonymous):

10x6.00=60 80x3.00=240

OpenStudy (phi):

yes, and 60+240 is 300

OpenStudy (anonymous):

240+60=300

OpenStudy (anonymous):

so ana sold her tickets for 300

OpenStudy (phi):

so your statement should say ana sold 10 senior tickets and 80 student tickets for a total of $300.00

OpenStudy (phi):

now if someone solves your problem they should get back 6 and 3 for the prices

OpenStudy (anonymous):

im supposed to solve my problem by using elimination substituition or graphing

OpenStudy (phi):

the good news is you know the answers

OpenStudy (phi):

Bridgette and ana are competing with each other to see who can sell the most tickets for their school play. Bridgette sold 20 senior tickets and 40 student tickets for a total of $240.00. Ana sold 10 senior tickets and 80 student tickets for a total of $300.00. Find the cost of one senior ticket and one student ticket.

OpenStudy (phi):

to solve, first should give names to the two unknowns. People usually use "x" and "y" what should the x stand for ? (either price of senior tickets or the price of student tickets)

OpenStudy (anonymous):

senior

OpenStudy (anonymous):

so x+y=240.00

OpenStudy (phi):

so you write: x is the price of a senior ticket y is the price of a student ticket

OpenStudy (phi):

you should think about what you wrote. You *know* x is 6.00 and y is 3.00 (you picked them). so x+y is not going to be 240.00 but you are on the right track. How do we get 240 ?

OpenStudy (anonymous):

20 * 6=120 40*3=120

OpenStudy (phi):

yes. but we are not supposed to know the answer. instead of using the number 6 we use the "variable" x (which will be 6 after we figure it out) in other words 20x is the total amount for the senior tickets (that means 20 times the (unknown) price)

OpenStudy (anonymous):

20x+40y=240.00

OpenStudy (phi):

exactly

OpenStudy (phi):

and you need a 2nd equation (using Ana's info)

OpenStudy (anonymous):

10x+80y=300.00

OpenStudy (phi):

20x+40y=240.00 10x+80y=300.00 can you solve this ?

OpenStudy (anonymous):

for elimination i could multiply 10 by -2 to eliminate the x's

OpenStudy (phi):

yes, so multiply the entire equation (both sides and all terms) by -2

OpenStudy (anonymous):

yea i know how to do the elimination process but how will i do the graphing how would i graph this

OpenStudy (phi):

I would simplify both equations by multiplying them by 1/10 that gives us 2x+4y=24 x+8y= 30 then I would multiply the 1st equation by ½ so we have x+2y=12 x+8y= 30 next, we plot these two lines. there are a few ways. one way is write them in slope-intercept form but all we need is 2 points, so I would pick x=0 and get 2y=12, y=6 or (0,6) is one point on the first line. then pick y=0, and we get x=12 so (12,0) is on the line. plot those two points, and "connect the dots" do the same thing for the other line. or use geogebra or some on-line tool, and automatically plot them

OpenStudy (anonymous):

thanks @phi